The one-time fling! Have you ever purchased an article of clothing (dress, sports jacket, etc.), worn the item once to a party, and then returned the purchase? This is called a one-time fling. About 15% of all adults deliberately do a one-time fling and feel no guilt about it! In a group of seven adult friends, find the following probabilities.. (a) No one has done a one-time fling. (Use 3 decimal places.). (b) At least one person has done a one-time fling. (Use 3 decimal places.). . ( c) No more than two people have done a one-time fling. (Use 3 decimal places.)

Verified Answer :

Using the binomial probability formula

b(x; n, p) = p^x nCx (1-P)^(n-x)

Where: 
x = flings 
n = sample size 

A) 
b(0; 8, .15)
= (0.15)^0 (8C0) (.85^8) 

= 1 1 .272 
= .272 

B) = 1 - A
= 1 - 0.272
= 0.728

C) 
b(1; 8, .15)
= (.15^1) (8C1) (.85^7 )

= .385 

b(2; 8, .15)
= (.15^2) (8C2) (.85^6) 

= .238 

b ≤ 2 = A + b1 + b2
= .272 + .385 + .238 
= .895

(a). The probability that no one has done one time fling is \(\boxed{\bf 0.377}\).

(b). The probability that at least one person has done one time fling is \(\boxed{\bf 0.623}\).

(c). The probability that no more than two person has done one time fling is \(\boxed{\bf 0.98279}\).

Further explanation:

Given:

If a person purchased an article of clothing and worn the cloths once then returned the clothes this is known as one time fling.

About \(15\%\) adults do one time fling.

Concept used:

The probability of an event \(E\) can be calculated as follows:

\(\boxed{P(E)=\dfrac{n(E)}{n(S)}}\)

Here, \(n(E)\) is the number of favorable outcomes in an event \(E\) and \(n(S)\) is the number of element in sample space \(S\).

The probability of exactly \(r\) success in \(n\) trial can be expressed as follows:

\(\boxed{P(F)=^{n}C_{r}p^{r}q^{n-r}}\)

Here, \(p\) is the probability of success in an event and \(q\) is the probability of failure.

Calculation:

Part (a):

The probability that adults do one time fling is \(0.15\).

The probability that no adult do one time fling is calculated as follows:

\(\boxed{1-0.15=0.85}\)

There are \(7\) adult friends in a group.

Consider \(A\) as an event that no one has done one time fling in a group of seven friends and \(P(A)\) as the probability of an event \(A\).

The probability \(P(A)\) can be calculated as follows:

\(\begin{aligned}P(A)&=^7C_{0}\cdot (0.15)^{0} \cdot (0.15)^{7-0}\\&=\dfrac{7!}{0!\cdot 7!}\cdot 1\cdot (0.85)^{7}\\&=1\cdot 1\cdot 0.377\\&=0.377\end{aligned}\)

Therefore, the probability \(P(A)\) is \(\boxed{\bf 0.377}\).

Part (b):

Consider \(A'\) as a complement event of an event \(A\).

The complement of event \(A\) is the event that at least one person has done one time fling in a group of \(7\) friends.

The probability of event \(A'\) can be calculated as follows:

\(\boxed{P(A')=1-P(A)}\)      …… (1)

Substitute \(P(A)=0.377\) in the equation (1) to obtain the probability   as follows:

\(\begin{aligned}P(A')&=1-0.377\\&=0.623\end{aligned}\)  

Therefore, the probability that at least one person has done one time fling in a group of \(7\) friends is \(\boxed{\bf 0.623}\).

Part (c):

The probability that no one has done one time fling is \(0.377\).

Consider \(B\) as an event that one person has done one time fling in a group of seven friends and \(P(B)\) as the probability of an event \(B\).

The probability \(P(B)\) can be calculated as follows:

\(\begin{aligned}P(B)&=^7C_{1}\cdot (0.15)^{1} \cdot (0.85)^{7-1}\\&=\dfrac{7!}{1!\cdot 6!}\cdot (0.15) \cdot (0.85)^{6}\\&=\dfrac{7!}{6!}\cdot 0.15 \cdot 0.37714\\&=7\cdot 0.0565\\&=0.396\end{aligned}\)

 

Consider \(C\) as an event that exactly two person has done one time fling in a group of seven friends and \(P(C)\) as the probability of an event \(C\).

The probability \(P(C)\) can be calculated as follows:

\(\begin{aligned}P(C)&=^7C_{2}\cdot (0.15)^{2} \cdot (0.85)^{7-2}\\&=\dfrac{7!}{2!\cdot 5!}\cdot (0.15)^{2}\cdot (0.85)^{5}\\&=\dfrac{7\cdot 6}{2}\cdot 0.0225\cdot 0.444\\&=21\cdot 0.00999\\&=0.20979\end{aligned}\)  

Now, the probability that no more than two person has done one time fling in a group of \(7\) friends can be calculated as follows:

\(\boxed{P(X)=P(A)+P(B)+P(C)}\)         …… (2)

Substitute \(P(A)=0.377\), \(P(B)=0.396\) and \(P(C)=0.209\) in the equation (1) to obtain the probability \(P(X)\) as follows:

\(\begin{aligned}P(X)&=0.377+0.396+0.20979\\&=0.98279\end{aligned}\)

Therefore, the probability that no more than two person has done one time fling is \(\boxed{\bf 0.98279}\).

Learn more:

1. Learn more about problem on numbers: https://brainly.com/question/1852063

2. Learn more about problem on function https://brainly.com/question/3225044

Answer details:

Grade: Senior school

Subject: Mathematics

Chapter: Probability

Keywords: Probability, exact event, sample space, number of element, complement event, success, failure, favorable, trial, one time fling, clothes, person, P(E)=n(E)/n(S).